The per-week (inverse) demand for use of the Øresund Bridge between Denmark and Sweden is P = 13 − 0.15Q during peak traffic periods and P = 10 − 0.1Q during off-peak hours, where Q is the number of cars crossing the bridge in thousands and P is the toll in euros. If the marginal cost of using the bridge is MC = 5 + 0.2Q, what is the optimal peak load toll and off-peak load toll for crossing the bridge?


1
Expert's answer
2021-05-24T13:11:34-0400

Demand during Peak "P = 13\u22120.15Q"

                "MC = 5+0.2Q"

Since,    "TR =P\u00d7Q"

 "TR =(13\u22120.15Q)Q"

          "TR = 13Q\u22120.15Q^2"

           "MR = \frac{∆TR}{∆Q}"

"MR = 13\u22120.30Q"


At optimal level     MC = MR

         "5+0.2Q\n\n= 13\u22120.30Q"

              "0.50Q = 8"

"Q=\frac{8}{0.50}"

"= 16"

"P =13\u22120.15(16)"

"=13\u22122.4 = 10.6"


Optimal peak load toll for crossing the bridge is 10.6

Demand during off-Peak load "P = 10\u22120.1Q"

"MC = 5+0.2Q"

Since,                "TR =P\u00d7Q"

 

                     "TR =(10\u22120.1Q)Q"

"TR = 10Q\u22120.1Q^\n\n2"

"MR=\frac{\Delta TR}{\Delta Q}"

"=10-0.20Q"


At optimal level,       MC = MR

               "5+0.2Q = 10\u22120.20Q"

              "0.40Q = 5\\Q=\frac{5}{0.40}=12.5"


"P =10\u22120.1(12.5) =10\u22121.25 =8.75"


Optimal off-peak load toll for crossing the bridge is 8.75